Note: (show all your works and measure that you answer all questions from part A to part B below. Sorry for long questions, but I give thanks to anyone who can help me)a. What is the pH of a buffer made by adding 10.00 mL of a 3.2 M HCOOH (aq) to 15.00 mL of a 3.2 M NaCOOH (aq) solution? pKa 3.74 b. What is the pH of a buffer made by dissolving 3.2 g of sodium acetate in 30 M acetic acid? The final volume is 1.0 L. Ka 1.8 x 10
Answer #1

a) pH = pKa + log [NaCOOH] / [HCOOH]

pKa = 3.74

[NaCOOH] = 15 x 3.2 / 30 = 1.6 M

[HCOoH] = 15 x 3.2 / 30 = 1.6 M

salt = acid so pH = pKa

pH = 3.74

b) [sodium acetate] = (3.2 / 82.03) (1 / 1) = 0.039 M

pH = pKa + log [sodium acetate] / [acetic acid]

pKa = - log [1.8 x 10-5] = 4.74

pH = 4.74 + log [0.039]/ [0.30]

pH = 4.74 - 0.89

pH = 3.85